# Haque, Dasgupta, and Rahman

### v1

@@ 4.1 Map Preprocessing and Grid Generation
At the beginning, we take an existing floorplan image of any valid format (e.g., PNG, JPEG, TIFF). We utilize the image to create a grayscale mask for binarization. This grayscale binary mask needs to be generated prior to running A* search and needs to be created only once for a specific map and can be reused for every subsequent iteration of the A* search. Figure 2 shows the entire steps of map image preprocessing for the floor plan of Orio Center, a shopping mall in Bergamo, Italy. The mask encodes walkable versus obstructed regions in the map. We load this mask and downsample it to a pre-determined $m\times n$ grid to produce a Boolean occupancy matrix $G$. The values of $m$ and $n$ depend on the aspect ratios of the map.
$G\in\{0,1\}^{m\times n},$ (1)
@@ 4.2 Graph Encoding
$c(\Delta i,\Delta j)=\{^{1,\text{ }|\Delta i|+|\Delta j|=1\text{ (orthogonal move)}}_{\sqrt{2},\text{ }|\Delta i|+|\Delta j|=1\text{ (diagonal move)}}$ (3)
@@
Table 1: Movement offsets and their associated costs in an eight-connected grid for A* search.
@@ 4.3 A* Path Planning with Diagonal Moves
A* is a classical best-first search algorithm that combines the exhaustive optimality of Dijkstra’s method with heuristic guidance to dramatically reduce the number of nodes required to examine for planning an optimal path [31]. This search algorithm is both (i) complete: it will surely find a path if one exists, (ii) optimal: it will return the most cost-optimal, given that the heuristic is admissible and consistent. In our grid-based search formulation, A* search maintains three quantities for each free cell $(i,j)$:
@@
    $g=g(i,j)+c((i,j)\rightarrow(i^{\prime},j^{\prime})),$ (4)
@@ 41  Prepend nodenode to pathpath;
Algorithm 1 A* Path Planning with 8-Way Connectivity
@@ 4.4 Path Compression
  • Stage 3 - Diagonal Collapse: Zig-zag walking patterns are unrealistic and need to be avoided. To address that, we scanned the run-length list for adjacent single-step pairs such as $(S,1)$ followed by $(E,1)$. Whenever found, we replaced those with a single diagonal entry such as $(SE,1)$. We then performed a second pass of run-length encoding to merge any resulting consecutive diagonal moves (e.g., combining two $(SE,1)$ into $(SE,2)$). A sample representation of the diagonally collapsed path and comparison with only the orthogonal path are shown in Figure 4(a) and 4(b).
@@
$(^{\prime}SE^{\prime},5),(^{\prime}N^{\prime},3),(^{\prime}E^{\prime},2)]$ (10)
@@ 5 Evaluation
We designed an experimental framework to evaluate each of the five pipeline stages. In the following sections, we describe the setup and measurement procedures for map preprocessing and grid generation, graph encoding, 8-way A* search, path compression, and language-model-based instruction synthesis. The average runtime of our approach and navigation instruction accuracies are also reported in this section. We conducted experiments on a desktop computer. Table 2 provides the configuration of the desktop computer. To be noted, all the model inferences and processings are done specifically using the CPU.
@@ 5.1 Map Preprocessing and Grid Generation
For our experimentation, we considered publicly available existing indoor maps of four different places: (i) Birmingham-Shuttlesworth International Airport, Alabama, USA. (ii) Milan Bergamo Airport, Bergamo, Italy; (iii) Bologna Guglielmo Marconi Airport, Bologna, Italy; and (iv) Orio Center, Bergamo, Italy. High-quality map images were collected for all these indoor locations. For grid generation, we utilized a Python script to load the map images as grid overlays of specific dimensions for each map. The grid dimensions were chosen based on the aspect ratios of the collected map images. Table 3 shows grid dimensions for each map images used in experimentation. An interactive grid layer was created over the map overlays to mark the walkable paths and blocked regions on the grid. We load an 8-bit grayscale image and partition it into the same $H\times W$ cell grid used during interactive editing. For each cell, we examine the corresponding pixel block in the map overlay: if more than half of its pixels are blocked by any obstacles in the actual map, we mark the cell free (1); otherwise, we mark it blocked (0). This majority-voting strategy handles noisy edges or slight misalignments in the hand-drawn mask, and guarantees that only cells designated as walkable remain passable. An example case of this is presented in Figure 5(a) and 5(b).
@@
Orio Center $1281:809$ $80\times 130$
@@ 5.2 Graph Encoding
After converting each map image into a regular grid mask, we construct a Boolean occupancy matrix in Python: free (walkable) cells are encoded as 1 and blocked (non-walkable) cells as 0. Behind the scenes, we leveraged NumPy’s slicing and vectorized comparisons to build this matrix. A sample $11\times 12$ occupancy matrix is shown in Figure 6.
@@ 5.3 Running A* Search with Path Compression
Orio Center $4.23\,\text{ms}$
The relatively simple layouts of the Birmingham and Bergamo maps translate into shorter average search times, whereas the larger, more intricate floorplans of Bologna and Orio Center naturally require more processing. Because A* is a fully deterministic graph search, it produces the exact same optimal path on every run, eliminating the variability of solely relying on a language model for wayfinding, and it computes these routes in just a few milliseconds, making it ideal for real-time indoor navigation.
@@ 5.4 SLM based Instruction Synthesis
For our experimentation, we used TinyLlama-1.1B, a small local language model with approximately 1.1 billion parameters, optimized for resource-constrained environments. This model supports multi-turn instruction-following and conversational dialogue generation, making it well-suited for natural language generation tasks [37]. The variant used in our experimentation has been fine-tuned using parameter-efficient fine-tuning (PEFT) techniques, specifically Low-Rank Adaptation (LoRA). This allowed us to specialize the model for this particular task while keeping memory and compute requirements reasonable. We trained the model on 1000 sample data points of converting terse commands into conversational output. All the inferences in our experimentation were performed using the CPU. Table 5 presents the average time in seconds required to run each query (A* Search + SLM) for different maps. It is evident that the whole process of A* pathfinding and SLM output generation takes 14-15 seconds on average.
@@
You are a precise navigation assistant. Convert provided terse commands into a numbered walking guide. Use “Start by walking…” for step 1, “Then walk…” for steps 2…(n-1), and “Finally walk…” for the last step. Output one numbered line per command.
Here is an example case given: Terse commands: Go East 3 steps Go North 1 step
For the example terse commands, the output is: 1. Start by walking east for 3 steps. 2. Finally, walk north for 1 step and you will reach your destination.
@@
Orio Center $16.96\,\text{s}$
Average $15.88\text{ s}$
@@ 5.5 Evaluation Outcomes
To assess the effectiveness of our proposed indoor navigation pipeline, we conducted a comparative evaluation against a closely related method introduced by Coffrini et al. [11], which serves as our baseline. Their framework represents one of the most comparable approaches in terms of objectives and application context, making it an appropriate reference point for benchmarking performance.
@@ 5.5.1 Baseline Method
Coffrini et al. [11] proposed an approach for generating indoor navigation instructions using LLM, specifically ChatGPT. In their implementation, the raw floor plan images of indoor circuits were preprocessed to remove any unnecessary information (texts, legends, etc.). They processed this image further by removing any extraneous visual elements as well, converting it to a simple graph-like representation of the connected components. This image, along with an iteratively refined system prompt and a user prompt specifying the source and destination, was provided as input to the LLM to generate human-readable navigation instructions.
While the method demonstrated a single-instruction-level accuracy of around 86%, it suffers from significant computational overhead. Specifically, the response time for generating a complete set of navigation instructions remains high, with an average processing time of approximately 4-5 minutes per query, mainly due to the image processing and reasoning capabilities required by the LLM [11].
@@ 5.5.2 Our Approach:
We conducted a total of 80 route-finding experiments, evenly distributed across the four test sites. Because our A* search is fully deterministic, it produced the exact same optimal path, and hence the exact same navigation instructions on every trial. In contrast, the purely LLM-based approach of Coffrini et al. often fails to match the intended floorplan geometry due to the model’s probabilistic nature, leading to incorrect or incomplete guidance. A comparison is shown in Table 6.
@@
Map Coffrini et al. Accuracy [11] Our Accuracy
@@
Orio Center 62.50% 100%
In their evaluation, Coffrini et al. [11] reported an accuracy of only 76% on the Bergamo floor plan, with performance declining to 62% on the larger and complex Orio Center layout. In contrast, our approach integrates A* pathfinding with an SLM, leveraging the A* algorithm for accurate path computation while reserving the machine learning model solely for the generation of natural language instructions. This modular design decouples the image processing and path planning from the language generation, ensuring the reasoning over spatial structures is handled algorithmically rather than probabilistically. As a result, our system achieved 100% accuracy across all experimented maps. Moreover, the reliance on lightweight local inference substantially reduced computational overhead. The average time required by A* to compute the optimal path is under 5 milliseconds, as shown in Table 4, and the total average time per iteration, including both pathfinding and SLM text generation, remains under 20 seconds (Table 5). This demonstrates practical usability in real-world scenarios. A comparison of processing times is shown in Figure 8.
@@
Taken together, these results confirm that by offloading deterministic path planning to A* and limiting the machine learning model’s role to lightweight text formatting, we can deliver both 100% route accuracy and true real-time performance, something that purely LLM-centric pipelines cannot match on resource-constrained devices.

### v2

@@ 3.1 Map Preprocessing and Grid Generation
At the beginning, we take an existing floorplan image of any valid format (e.g., PNG, JPEG, TIFF). Any unnecessary parts (i.e., legends, annotations, etc.) that would not be helpful for navigation from the map image are cropped and removed. We utilize the map image to create a grayscale mask for binarization. This grayscale binary mask needs to be generated prior to running A* search and needs to be created only once for a specific map, and can be reused for every subsequent iteration of the A* search. Figure 3 shows the entire steps of map image preprocessing for the floor plan of Orio Center, a shopping mall in Bergamo, Italy.
Figure 2: Portal nodes at the connected points between the ground floor and first floor of Orio Center, Bergamo, Italy.
The mask encodes walkable versus obstructed regions in the map. We load this mask and downsample it to a pre-determined $m\times n$ grid to produce a Boolean occupancy matrix $G$. For locations comprising multiple floors, a separate occupancy matrix is constructed for each level. To enable vertical traversal, portal nodes are manually defined at the locations of the connecting points (i.e., staircases, elevators, or escalators). Figure 2 shows an example case over two floors of Orio Center, Bergamo, Italy. A portal is represented as a bidirectional link between two grid cells belonging to different floors, formally encoded as an edge in the navigation graph coordinates $(f_{1},i_{1},j_{1})$ on floor $f_{1}$ to coordinates $(f_{2},i_{2},j_{2})$ on floor $f_{2}$. The values of $m$ and $n$ for the Boolean occupancy matrix depend on the aspect ratios of the map.
$G\in\{0,1\}^{m\times n},$
@@ 3.2 Graph Encoding
$c(\Delta i,\Delta j)=\{^{1,\text{ }|\Delta i|+|\Delta j|=1\text{ (orthogonal move)}}_{\sqrt{2},\text{ }|\Delta i|+|\Delta j|=1\text{ (diagonal move)}}$
@@
Table 1: Movement offsets and their associated costs in an eight‐connected grid for A* search.
@@ 3.3 A* Path Planning with Diagonal Moves
A* is a classical best-first search algorithm that combines the exhaustive optimality of Dijkstra’s method with heuristic guidance to dramatically reduce the number of nodes required to examine for planning an optimal path Hart et al., (1968). This search algorithm is both (i) complete: it will surely find a path if one exists, (ii) optimal: it will return the most cost-optimal, given that the heuristic is admissible and consistent. In our grid-based search formulation, A* search maintains three quantities for each free cell $(i,j)$:
@@
    $g=g(i,j)+c((i,j)\rightarrow(i^{\prime},j^{\prime})),$
@@ 41  Prepend nodenode to pathpath;
Algorithm 1 A* Path Planning with 8‑Way Connectivity
@@ 3.4 Path Compression
  • Stage 3 - Diagonal Collapse: Zig-zag walking patterns are unrealistic and need to be avoided. To address that, we scanned the run-length list for adjacent single-step pairs such as $(S,1)$ followed by $(E,1)$. Whenever found, we replaced those with a single diagonal entry such as $(SE,1)$. We then performed a second pass of run-length encoding to merge any resulting consecutive diagonal moves (e.g., combining two $(SE,1)$ into $(SE,2)$). A sample representation of the diagonally collapsed path and comparison with only the orthogonal path are shown in Figure 5(a) and 5(b).
@@
$(^{\prime}SE^{\prime},5),(^{\prime}N^{\prime},3),(^{\prime}E^{\prime},2)]$
@@ 4 Evaluation
We designed an experimental framework to evaluate each of the five pipeline stages. In the following sections, we describe the setup and measurement procedures for map preprocessing and grid generation, graph encoding, 8‑way A* search, path compression, and language‑model‑based instruction synthesis. The average runtime of our approach and navigation instruction accuracies are also reported in this section. We conducted experiments on a desktop computer. Table 2 provides the configuration of the desktop computer. To be noted, all the model inferences and processings are done specifically using the CPU.
@@ 4.1 Map Preprocessing and Grid Generation
For our experimentation, we considered publicly available existing indoor maps of four different places: (i) Birmingham-Shuttlesworth International Airport, Alabama, USA. (ii) Milan Bergamo Airport, Bergamo, Italy; (iii) Bologna Guglielmo Marconi Airport, Bologna, Italy; and (iv) Orio Center, Bergamo, Italy. High-quality map images were collected for all these indoor locations. For grid generation, we utilized a Python script to load the map images as grid overlays of specific dimensions for each map. The grid dimensions were chosen based on the aspect ratios of the collected map images. Table 3 shows grid dimensions for each map images used in experimentation. An interactive grid layer was created over the map overlays to mark the walkable paths and blocked regions on the grid. We load an 8‑bit grayscale image and partition it into the same $H\times W$ cell grid used during interactive editing. For each cell, we examine the corresponding pixel block in the map overlay: if more than half of its pixels are blocked by any obstacles in the actual map, we mark the cell blocked (0); otherwise, we mark it free (1). This majority‑voting strategy handles noisy edges or slight misalignments in the hand‑drawn mask, and guarantees that only cells designated as walkable remain passable. An example case of this is presented in Figure 6(a) and 6(b).
@@
Orio Center (Ground Floor) $1281:809$ $80\times 130$
Orio Center (First Floor) $1169:947$ $80\times 100$
@@ 4.2 Graph Encoding
After converting each map image into a regular grid mask, we construct a Boolean occupancy matrix in Python: free (walkable) cells are encoded as 1 and blocked (non‑walkable) cells as 0. Behind the scenes, we leveraged NumPy’s slicing and vectorized comparisons to build this matrix. A sample $11\times 12$ occupancy matrix is shown in Figure 7.
@@ 4.3 Running A* Search with Path Compression
Orio Center (Ground Floor) $4.23\,\text{ms}$
Orio Center (First Floor) $4.09\,\text{ms}$
Orio Center (Multi-floor) $8.80\,\text{ms}$
The relatively simple layouts of the Birmingham and Bergamo maps translate into shorter average search times, whereas the larger, more intricate floor plans of Bologna and Orio Center naturally require more processing. Because A* is a fully deterministic graph search, it produces the exact same optimal path on every run, eliminating the variability of solely relying on a language model for wayfinding, and it computes these routes in just a few milliseconds, making it ideal for real‑time indoor navigation.
@@ 4.4 SLM based Instruction Synthesis
For our experimentation, we used TinyLlama-1.1B, a small local language model with approximately 1.1 billion parameters, optimized for resource-constrained environments. This model supports multi-turn instruction-following and conversational dialogue generation, making it well-suited for natural language generation tasks Zhang et al., (2024). The variant used in our experimentation has been fine-tuned using parameter-efficient fine-tuning (PEFT) techniques, specifically Low-Rank Adaptation (LoRA). This allowed us to specialize the model for this particular task while keeping memory and compute requirements reasonable. We trained the model on 1000 sample data points of converting terse commands into conversational output. All the inferences in our experimentation were performed using the CPU. Table 5 presents the average time in seconds required to run each query (A* Search + SLM) for different maps. It is evident that the whole process of A* pathfinding and SLM output generation takes 14-15 seconds on average.
@@
You are a precise navigation assistant. Convert provided terse commands into a numbered walking guide. Use “Start by walking…” for step 1, “Then walk…” for steps 2…(n-1), ”Take the escalator from Floor n to n+1”, and “Finally walk…” for the last step. Output one numbered line per command.
Here is an example case given: Terse commands: Go East 3 steps Take the escalator from Floor 0 to 1 Go North 1 step
For the example terse commands, the output is: 1. Start by walking east for 3 steps. 2. Take the escalator from Floor 0 to 1. 2. Finally, walk north for 1 step, and you will reach your destination.
@@
Orio Center (Ground Floor) $16.96\,\text{s}$
Orio Center (Multi-floor) $21.09\,\text{s}$
Average $16.92\,\text{s}$
@@ 4.5 Evaluation Outcomes
To assess the effectiveness of our proposed indoor navigation pipeline, we conducted a comparative evaluation against a closely related method introduced by Coffrini et al. Coffrini et al., (2025), which serves as our baseline. Their framework represents one of the most comparable approaches in terms of objectives and application context, making it an appropriate reference point for benchmarking performance.
@@ 4.5.1 Baseline Method
Coffrini et al. Coffrini et al., (2025) proposed an approach for generating indoor navigation instructions using LLM, specifically ChatGPT. In their implementation, the raw floor plan images of indoor circuits were preprocessed to remove any unnecessary information (texts, legends, etc.). They processed this image further by removing any extraneous visual elements as well, converting it to a simple graph-like representation of the connected components. This image, along with an iteratively refined system prompt and a user prompt specifying the source and destination, was provided as input to the LLM to generate human-readable navigation instructions.
While the method demonstrated a single-instruction-level accuracy of around 86%, it suffers from significant computational overhead. Specifically, the response time for generating a complete set of navigation instructions remains high, with an average processing time of approximately 4-5 minutes per query, mainly due to the image processing and reasoning capabilities required by the LLM Coffrini et al., (2025).
@@ 4.5.2 Our Approach:
We conducted a total of 80 route-finding experiments, evenly distributed across the four test sites. Because our A* search is fully deterministic, it produced the exact same optimal path, and hence the exact same navigation instructions on every trial. In contrast, the purely LLM‐based approach of Coffrini et al. often fails to match the intended floorplan geometry due to the model’s probabilistic nature, leading to incorrect or incomplete guidance. A comparison is shown in Table 6.
@@
Map Coffrini et al. Accuracy Coffrini et al., (2025) Our Accuracy
@@
Orio Center (Ground Floor) 62.50% 100%
Orio Center (Multi-floor) N/A 100%
In their evaluation, Coffrini et al. Coffrini et al., (2025) reported an accuracy of only 76% on the Bergamo floor plan, with performance declining to 62% on the larger and complex Orio Center layout. In contrast, our approach integrates A* pathfinding with an SLM, leveraging the A* algorithm for accurate path computation while reserving the machine learning model solely for the generation of natural language instructions. This modular design decouples the image processing and path planning from the language generation, ensuring the reasoning over spatial structures is handled algorithmically rather than probabilistically. As a result, our system achieved 100% accuracy across all experimented maps. Moreover, the reliance on lightweight local inference substantially reduced computational overhead. The average time required by A* to compute the optimal path is under 5 milliseconds, as shown in Table 4, and the total average time per iteration, including both pathfinding and SLM text generation, remains under 20 seconds (Table 5). This demonstrates practical usability in real-world scenarios. A comparison of processing times is shown in Figure 9.
@@
Taken together, these results confirm that by offloading deterministic path planning to A* and limiting the machine learning model’s role to lightweight text formatting, we can deliver both 100% route accuracy and true real‑time performance, something that purely LLM‑centric pipelines cannot match on resource‑constrained devices.